OpenStudy (kaitlyn_nicole):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2.

OpenStudy (aravindg):

F(g(x)) means function f operates on g. So replace the x in -x+4 by expression for g(x). Simplify that anr then plug in x=2 to get your answer.

OpenStudy (denonakavro):

So first plug in x in the equation of g -2(2) -3 -4-3= -7 Now put the -7 in -x +4 in the place of x -(-7)+4 7+4 11 will be your answer

OpenStudy (dayakar):

f(x) =-x +4 -------(1) g(x) = -2x -3-----(2) f(g(x) = f(-2x-3) [ from (2) ] =-(-2x-3) +4 [ from (1) ] = 2x+3+4 = 2x+7 substitute x= 2 u can get the answer